OpenStudy (james1769):

Solve the equation the square root of the quantity x + 4 - 3 equals 1 for the variable. Show each step of your solution process.

OpenStudy (michele_laino):

is your equation like this? \[\sqrt {x + 4} - 3 = 1\]

OpenStudy (james1769):

yes

OpenStudy (michele_laino):

since we have this square root \(\sqrt {x + 4} \), then we have to request that the radical \(x+4\) has to be greater or equal than zero: \[x + 4 \geqslant 0\]

OpenStudy (michele_laino):

since the square root of negative numbers doesn't exists

OpenStudy (james1769):

okay thanks

OpenStudy (michele_laino):

solving that inequality we have: \[x \geqslant - 4\] which means that we can accept, as solutions, only those quantities which are greater or equal to -4

OpenStudy (james1769):

does the answer come out to be x=12

OpenStudy (james1769):

got one more question

OpenStudy (michele_laino):

yes! since we have to add 3 to both sides: \[\sqrt {x + 4} = 4\] then squaring both sides and simplifying, we get: x=12

OpenStudy (michele_laino):

and since 12 > -4, we can conclude that x=12 is the solution of our equation

OpenStudy (james1769):

one more question

OpenStudy (michele_laino):

ok!